I don't understand how this problem get's set up
The supplemental angle added to the complement of the angle gives 210 degrees. What is the measure of the angle?

Found 2 solutions by mananth, ankor@dixie-net.com:
Answer by mananth(16946)

let angle be x
...
supplement = 180-x
complement = 90-x
..
180-x+90-x=210
-2x+270=210
-270
-2x=-270+210
-2x=-60
/-2
x=30 the angle

Answer by ankor@dixie-net.com(22740)

The supplemental angle added to the complement of the angle gives 210 degrees.
What is the measure of the angle?
:
Let a = the angle
then we know
(180-a) = the supplement of the angle
and
(90-a) = the complement of the angle
:
Therefore the equation is:
(180-a) + (90-a) = 210
-2a + 270 = 210
-2a = 210 - 270
-2a = -60
a = 30 degrees is the angle, not that hard, right?